OpenStudy (anonymous):

Using the subsitution y = log(base 3) x, solve the below equation

OpenStudy (anonymous):

\[\log_{3} x + 2 = 3 \log_{x}3\]

OpenStudy (anonymous):

you can start by converting the log to base 10 by going log(x)/log(3) and the other one to log(3)/log(x)

OpenStudy (anonymous):

from there all u have to do is isolate for x. can u do that?

OpenStudy (anonymous):

Ok, I have the answer thanks for your help

terenzreignz (terenzreignz):

I wouldn't do that if I were you... just this use this property \[\huge \log_b a = \frac1{\log_a b}\]

terenzreignz (terenzreignz):

So that the equation just becomes \[\huge \log_3x+2=\frac3{\log_3 x}\] Now substitute the y, as instructed \[\huge y + 2 = \frac3y\]And isn't that much better? :)